OpenStudy (anonymous):

how do i integrate sqrt(1+x^3)??

OpenStudy (anonymous):

By part integral may be :\

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

just try it :P

OpenStudy (lgbasallote):

someone once told me this cannot be solved by usual means

OpenStudy (lgbasallote):

you'll have to use a method like "numerical analysis" or something

OpenStudy (anonymous):

by parts is promising..

OpenStudy (lgbasallote):

http://www.wolframalpha.com/input/?i=int+%28sqrt+%281%2Bx%5E3%29%29dx <--cannot be solved by usual means

OpenStudy (anonymous):

"by parts" - @Pinguine it is NOT a funny joke. Just childish cruelty to the less informed...

OpenStudy (anonymous):

I just try to do what I used to do, if it cannot be solved so It can't. you're clever than me so give him an advice. sir

OpenStudy (anonymous):

By parts - WHEN THERE ARE ANY PARTS, Don't you think ?!!

OpenStudy (anonymous):

By the way, definitely the worst thing people experience as tourists when they are given wrong answers, by someone who knows in advance that he has no real good answer.

OpenStudy (anonymous):

I just tell him that's "it may be". If I make you confuse so I would apologize, OK ??

OpenStudy (anonymous):

Sorry too

OpenStudy (anonymous):

itergrate fr (1+x^3)^(1/2) = 2(1+x^3)^(1/2+2/2) = 2(1+x^3)^(3/2) isnt it? haaah. cmiiw

OpenStudy (anonymous):

is it x to the power 3?